OpenStudy (anonymous):

Consider the line that passes through the point and is parallel to the given vector. (1, -5, 9) ‹-1, 2, -3› (a) Find symmetric equations for the line.

OpenStudy (anonymous):

x= -1t+1 y=2t-5 z= -3t+9

OpenStudy (anonymous):

this is parametric, we want symmetric,

OpenStudy (anonymous):

x-1=-t -x+1=t

OpenStudy (anonymous):

y+5=2t (y+5)/2 =t

OpenStudy (anonymous):

so, how do you get parametric eq's out of that? and then symmetric is just parametric solved for t?

OpenStudy (anonymous):

yes, solve for t to get symmetric

OpenStudy (anonymous):

Example ------------ vector <a,b,c> point <p,q,r> x= at +p y= bt+q z= ct+ r

OpenStudy (anonymous):

awesome.

OpenStudy (anonymous):

so t=(z-9)/-3

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

yep, a line is just a vector add to a point.

OpenStudy (amistre64):

since the vector can be any scaled version of itself; we use a "t" to indicate that any size of this vector is sufficient to produce a line